tennis balls are sold in packages of three balls.
number of balls / number of packages
30 / 10
72 / 24
378 / 126
999 / 333
t /
I can't remember how to find t. I don' see a pattern.

Answer by checkley77(12844)

t=30/10=3 tennis balls per package.
t/(1/3)t
t/(t/3)
t*3/t=3t/t=3 ans.